# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

### v1

@@ 2.7 Wage bargaining
In each period, matched worker-firm pairs bargain for wages over the matched surplus. A worker has their matched surplus as $(W_{i}(R,s,a)-U^{I}_{i}(R,s,a))$ and a firm has its matched surplus as $J_{i}(Ra,s,a)$, where $i$ denotes $np$ or $p$. The worker’s bargaining power $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$. The resulting wage for workers at $p$ and $np$ firms are given by Equation 20 and Equation 21, respectively.
@@
We define the stationary equilibrium in Appendix A.
@@ 3 Calibration
A period in the model represents a quarter. We calibrate the benchmark model to match the relevant U.S. economy moments. We have two sets of parameters. The first group of parameters is chosen externally based on literature and empirical evidence without using model-generated data, while the second group of parameters is calibrated internally by simulating our model to match a set of relevant data moments. We set the share of black workers to be 19.1% in the model, with the remaining being white workers. Table 1 shows the internally calibrated parameter values and their targeted moments, as well as the externally chosen parameters and their sources.
@@
$\kappa_{\text{np}}$ 2.7350 np sector vacancy posting cost market tightness 1 1
@@
$\alpha$ 0.6600 elasticity of labor matching Nakajima (2012)
$\theta_{n}$ 0.2890 capital share of output Nakajima (2012)
@@
$\xi$ 0.8433 maximum UI coverage Setty and Yedid-Levi (2021) 48% median wage
@@
Notes: This table reports the parameters, their values, and descriptions. The top panel presents the parameters chosen internally by minimizing the distance between model generated moments and data. The last two columns of the top panel compare the targeted moments in the data and in the model simulations. The bottom panel reports the parameters chosen externally, their values, and description.
@@ 3.2 Production
The worker-firm match produces according to a Cobb-Douglas production function, $f(k)=k^{\alpha}$. We choose $\alpha$ to be 0.289 and set the quarterly capital depreciation rate $\delta$ equal to 0.015 following Nakajima (2012).
@@ 3.3 Productivity and wealth shocks
We use hourly real wage from the Panel Study of Income Dynamics (PSID) to estimate the persistence, $\rho_{s}$, and the standard deviation, $\sigma_{s}$ of the productivity process. Our estimation strategy closely follows Setty and Yedid-Levi (2021). We run a standard Mincer wage regression of log wage on the demographic controls, including education, labor market experience, race, marital status, year, and state fixed effects for a sample of males aged 25 years and above. We then use the obtained residuals to estimate an AR(1) regression by utilizing the panel dimension of the PSID data. The estimated regression is at a biennial frequency, given that the PSID data is available once in two years. We assign the quarterly adjusted coefficient of the AR(1) regression as the persistence of the idiosyncratic productivity process. The standard deviation of the productivity process corresponds to the standard deviation of the residuals from the AR(1) regression after adjusting for the model frequency. We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, as equal to 0.1680.
Following Mukoyama (2010), our model also features extreme wealth shocks to capture the mass of people having zero wealth. The race-specific probability of losing one’s wealth, $\epsilon_{R}$, is calibrated to be 0.0179 for black workers and 0.0086 for white workers. They capture the empirical moments that around 18% of black workers and 7% white workers have zero wealth. (Nakajima, 2021).
@@ 3.4 Unemployment insurance
The unemployment insurance system in our model is characterized by the replacement rate, $h$, maximum insurance payout, $\chi$, and the probability of maintaining the eligibility status, $P_{e}$. We choose the replacement rate, $h$, to be 0.4 along the lines of Shimer (2005), Mitman and Rabinovich (2015), and others. Following Setty and Yedid-Levi (2021), we calibrate the maximum payout $\chi$ to be 0.8433, which amounts to 48% of the median wage in the model. The eligibility probability $P_{e}$ is chosen to be 0.5385 to generate an average unemployment benefit duration of 26 weeks, as in Mitman and Rabinovich (2015).
@@ 3.5 Labor search and wage bargain
The labor market turnover statistics are obtained from Cajner et al. (2017). The elasticity of the matching function, $\iota$, targets the job finding rate of the black workers, and our calibrated value of 1.3012 is close to that of 1.25 in Den Haan et al. (2000) and Petrosky-Nadeau et al. (2018). We choose firm-type specific job destruction shocks $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers respectively. The ensuing values for $\lambda_{np}$ and $\lambda_{p}$ are 0.0644 and 0.0268, respectively. [^5] All the black workers work only with $np$ firms, and hence $\lambda_{np}$ is exactly equal to the job separation rate of the black workers. On the other hand, white workers work in both $p$ and $np$ firms, and hence $\lambda_{p}$ (0.0268) is smaller than the aggregate separation rate of the white workers (0.0380). The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job finding rate of white workers, while the posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to target the aggregate labor market tightness, $\theta$, to be 1 following Wolcott (2021). [^6] Aggregate labor market tightness $\theta=\frac{v_{np}+v_{p}}{u_{bl}+u_{wh}}$. We find that the prejudiced firms pay a lot more than non-prejudiced firms to post their vacancies, with $\kappa_{p}$ calibrated to be 4.2622, compared to $\kappa_{np}$ taking a value of 2.735. The bargaining power of black workers, $\xi_{bl}$, is calibrated to target the average black-white racial wage ratio of 0.75 (Derenoncourt and Montialoux, 2021). The white worker’s bargaining power, $\xi_{wh}$, is chosen to generate the average firm profit share of 3.3% as in Nakajima (2012). Consistent with our expectations, we find that the black workers have a lower bargaining power compared to the white workers, with $\xi_{bl}$ to be 0.1397 and $\xi_{wh}$ to be 0.211.
@@ 3.6 Assessing the model as a quantitative theory of racial disparity
Our calibration successfully captures the racial gaps along three important dimensions: income, wealth, and labor market outcomes. We now discuss each of these dimensions in more detail before proceeding with the quantitative exercises examining the macroeconomic impact of racial discrimination.
@@
Note: This figure compares the steady state model generated distributions of wealth and labor income with their empirical analogs. The empirical labor income and wealth distributions are estimated from the 2007 American Community Survey (ACS) and the 2007 Survey of Consumer Finance (SCF), respectively.
First, the model successfully captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we don’t target these distributions explicitly. As with the majority of incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile. On the other hand, we are much closer to the empirical distributions in the lower quintiles since we target the share of zero-wealth workers in our calibration.
@@
Second, the model is able to capture the disparities in unemployment, labor income, and wealth between black and white workers. As targeted in the calibration, the model replicates the empirical racial labor income gap and the empirical racial difference in the share of people at zero wealth. In addition to the targeted moments, Table 2 demonstrates that our model successfully captures the untargeted employment and wealth moments. Specifically, our model reproduces the 5% unemployment rate among white workers compared to 12% among black workers. Apart from the lower separation rate, white workers have a lower unemployment rate due to their access to the prejudiced sector. Our model results show that the $p$ sector with the equilibrium market tightness of 0.4, though smaller than the $np$ market tightness of 0.73, provides an important advantage for white workers by increasing their job-finding rate compared to black workers. On the wealth dimension, without targeting, the model generates a mean black-white wealth ratio of 0.28, close to 0.23 as in Kaplan et al. (2014). However, the model understates the median wealth gap between black and white workers. The median black worker holds 33% of the wealth of the median white worker in our model compared to 17% in the data (Kaplan et al., 2014).
In sum, the model captures the racial inequalities across labor income, wealth, and unemployment outcomes. We next explore how racial discrimination in hiring contributes to these differences.
@@ 4.1 Sustaining racial discrimination
The labor search-and-matching framework allows firms to earn profit from the surplus generated by the search frictions. This feature permits prejudiced and non-prejudiced firms to coexist in a competitive equilibrium. It contrasts with a frictionless canonical neoclassical model where prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers as described in Becker (1957). In particular, our benchmark model has $p$ firms retaining a profit of 0.04 while $np$ firms retain a profit of 0.03 in the steady state (see Column 1 of Table 3).
Next, we examine the steady-state interplay between discriminatory and non-discriminatory firms. Since discriminatory hiring remains profitable in the benchmark equilibrium, we gradually increase penalties by raising the vacancy posting costs for discriminatory firms until they are no longer profitable. Figure 2 displays the dynamics of employment, vacancy, and profits of the prejudiced and non-prejudiced firms. As $\kappa_{p}$ rises, $p$ firms become increasingly disadvantaged, gradually losing their profits, vacancy postings, and employment. Conversely, with fewer $p$ firms competing in the market, $np$ firms see a steady increase in profits, vacancy postings, and employment.
It is worth noting that the dynamics of $p$ and $np$ firms are not symmetric, given the differences in vacancy posting costs and job destruction shocks. As $p$ firms’ profit declines, the profit of $np$ firms increases at a slower rate. Additionally, as $np$ firms replace $p$ firms, they recruit from both black and white workers, rather than only white workers displaced from $p$ firms. This is mainly reflected in $np$ firms’ more elastic vacancy posting response.
In the limit, as $\kappa_{p}$ approaches the total expected value of the $p$ sector, discriminatory ($p$) firms almost completely exit the market, since posting vacancies exclusively for white workers becomes unprofitable. The value of $\kappa_{p}$ at which $p$ firms exit is 217% of the benchmark value of 4.26. Meanwhile, non-discriminatory ($np$) firms become nearly the sole employers in the economy. The gray dotted lines in Figure 2 illustrate a counterfactual steady state where we entirely remove $p$ firms from the benchmark model. Thus, increasing $\kappa_{p}$ in the limit asymptotically equals the alternative of completely eliminating $p$ firms.
@@ 4.2 Impact of hiring discrimination on racial inequality
In this subsection, we compare the benchmark economy to one without prejudiced firms to illustrate the aggregate and distributional impact of discriminatory hiring. Table 3 presents the comparisons. All firms in the counterfactual models provide equal hiring opportunities to black and white workers. Column 2 removes the $p$ firms from the model, which leaves all workers with equal job destruction shocks at the benchmark $\lambda_{np}$. In Column 3, we re-calibrate the job destruction shocks to race-specific and have $\lambda_{wh}$ and $\lambda_{bl}$ as race-specific job destruction shocks. [^7] In the benchmark economy, we structure firm-specific job destruction shocks ($\lambda_{np}$ and $\lambda_{p}$) to match race-specific job separation rates. This counterfactual economy only has one type of firm. We assign white workers a job destruction shock, $\lambda_{wh}$, to be 0.038 and black workers a destruction shock, $\lambda_{bl}$, to be 0.0644 to keep black and white workers’ job separation rates as in the calibration. Re-assigning the job destruction shocks further isolates the impact of job separation.
@@
The first panel shows the changes in racial employment, wage, and wealth disparities after removing prejudiced hiring. The first two rows describe the job separation rates. By construction, Column 2 equalizes job separation rates by race, while Column 3 retains the difference as the benchmark. The next two rows describe the job-finding rates. Since there is no $p$ sector in the economy in Columns 2 and 3, the job-finding rates are equalized for black and white workers. White workers only have one sector for employment. Hence, their job-finding rate reduces. Black workers have more job opportunities as the $np$ sector posts more vacancies; hence, their job-finding rate increases. Comparing Columns 2 and 3, the much lower job destruction shock for white workers in Column 3 results in fewer workers searching for jobs, which leads to higher job-finding rates than in Column 2. Altogether, it reduces the racial gap in the unemployment rate. In Column 2, the racial unemployment rate gap completely closes as job turnover equalizes. In Column 3, the unemployment rate gap closes from 7% to 3% because of the remaining racial difference in $\lambda$.
Turning to the wage impact, black workers can bargain for a higher wage rate as they receive more job opportunities when we remove discriminatory hiring. White workers, however, no longer have access to the exclusive $p$ sector; their outside options are lower during bargaining, thus driving down their wage rate. As a result, the average black-to-white wage ratio increases from 75% in benchmark to 88% in Column 2 and to 86% in Column 3. We conclude that removing discriminatory hiring closes 52% (Column 2) to 44% (Column 3) of the average wage gap between black and white workers.
Additionally, removing discriminatory hiring reduces the wealth gap, with the mean wealth ratio moving from 0.28 to 0.33 and the median wealth ratio increasing from 0.33 to 0.44. The result remains stable between Columns 2 and 3. We infer that discriminatory hiring explains about 7% of the mean wealth gap and 16% of the median wealth gap.
@@
The second panel displays the changes in firm attributes. We loosely refer to the one-sector economy without discriminatory hiring as $np$ firms, though the Column 3 specification differs from the benchmark $np$ firms in their job destruction shock. Without discriminatory firms, $np$ firms’ profit doubles, vacancy posting, and employment level increase. These changes result in the aforementioned advantages for black workers in the counterfactual economies. In Column 3, white workers’ lower job destruction shock drives up the overall $np$ employment and reduces vacancy postings compared to Column 2.
The third panel of Table 3 displays the labor market conditions. In response to the $p$ firms exiting, more $np$ firms endogenously enter the market. The labor market tightness of the $np$ sector, $\theta_{np}$, jumps from 0.73 in the benchmark to 1.02 in the equilibrium without $p$ firms and 1.12 when we split $\lambda$ by race. The ensuing market tightness is higher than the benchmark aggregate market tightness of 1.
The last panel of Table 3 describes the aggregate outcomes. With more $np$ firms entering, the aggregate output stays relatively unchanged in Column 3. The capital-output ratio slightly decreases to 10.25. The economy-wide average wage increases to 2.05 from benchmark 2.04, and the aggregate unemployment rate decreases to 6%. However, Column 2 experiences a more significant reduction in output, average wage, and higher unemployment rate due to the overall higher job destruction shock that requires firms to frequently pay for vacancy posting costs rather than a resource for production.
Our results show that the presence of discriminatory firms can explain sizable fractions of income and wealth inequality. Penalizing prejudiced firms primarily improves the employment opportunities and wage outcomes for the black workers; it also worsens the situation for the white workers. Removing $p$ firms reduces the job-finding rate and, hence, the unemployment rate for white workers. In addition, they also lower the value of their outside option in wage bargaining, which, in turn, affects their income and wealth outcomes. To quantify the differential effects of removing discriminatory hiring, we next calculate the welfare changes experienced by the black and the white workers.

### v2

@@ 2.7 Wage bargaining
In each period, matched worker-firm pairs bargain for wages over the matched surplus. A worker has their matched surplus as $(W_{i}(R,s,a)-U^{I}_{i}(R,s,a))$ and a firm has its matched surplus as $J_{i}(Ra,s,a)$, where $i$ denotes $np$ or $p$. A worker’s bargaining power $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$. The resulting wage for workers at $p$ and $np$ firms are given by Equation 20 and Equation 21, respectively.
@@
Finally, we define the stationary equilibrium in Appendix A.
@@ 3 Calibration
Each period in the model represents a quarter. We calibrate the benchmark model to match the relevant U.S. economy moments. We have two sets of parameters. One group of parameters is chosen externally based on literature and empirical evidence without using model-generated data. The other parameters are calibrated internally by simulating our model to match a set of relevant data moments. We set the share of black workers to be 19.1% in the model, with the remaining being white workers. Table 1 shows the internally calibrated parameter values, their targeted moments, the externally chosen parameters, and their sources.
@@
$\kappa_{\text{np}}$ 2.7350 np sector vacancy posting cost market tightness 1.0 1.0
@@
$\alpha_{k}$ 0.2890 capital share of output Nakajima (2012)
@@
$\chi$ 0.8433 maximum UI coverage Setty and Yedid-Levi (2021) 48% median wage
@@
Notes: This table reports the parameters, their values, and descriptions. The top panel presents the parameters chosen internally by minimizing the distance between model-generated moments and data. The last two columns of the top panel compare the targeted moments in the data and in the model simulations. The bottom panel reports the parameters chosen externally, their values, and description.
@@ 3.2 Production
The worker-firm match produces according to a Cobb-Douglas production function, $f(k)=k^{\alpha_{k}}$. We choose $\alpha_{k}$ to be 0.289 and set the quarterly capital depreciation rate $\delta$ equal to 0.015, following Nakajima (2012).
@@ 3.3 Productivity and wealth shocks
We use hourly real wage from the Panel Study of Income Dynamics (PSID) to estimate the persistence, $\rho_{s}$, and the standard deviation, $\sigma_{s}$ of the productivity process. Our estimation strategy closely follows Setty and Yedid-Levi (2021). We run a standard Mincer wage regression of log wages on the demographic controls, including education, labor market experience, race, marital status, year, and state fixed effects for a sample of men 25 years and older. We then use the obtained residuals to estimate an AR(1) regression by utilizing the panel dimension of the PSID data. The estimated regression is at a biennial frequency, given that the PSID data is available once every two years. We assign the quarterly adjusted coefficient of the AR(1) regression as the persistence of the idiosyncratic productivity process. The standard deviation of the productivity process corresponds to the standard deviation of the residuals from the AR(1) regression after adjusting for the model frequency. We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, as equal to 0.1680.555We follow Setty and Yedid-Levi (2021) to have only men in the sample. We also calculate the AR(1) processes, including men and women from PSID. The resulting $\rho_{s}$ is unchanged from the men-only sample (0.9411), and $\sigma_{s}$ is slightly lower (0.1671) than the men-only sample.
Following Mukoyama (2010), our model also features extreme wealth shocks to capture the mass of people having zero wealth. The race-specific probability of losing one’s wealth, $\epsilon_{R}$, is calibrated to be 0.0179 for black workers and 0.0086 for white workers. They capture the empirical moments that around 18% of black workers and 7% white workers have zero wealth (Nakajima, 2021).
@@ 3.4 Unemployment insurance
The unemployment insurance system in our model is characterized by the replacement rate, $h$, maximum insurance payout, $\chi$, and the probability of maintaining the eligibility status, $P_{e}$. We choose the unemployment insurance replacement rate, $h$, to be 0.4 along the lines of Shimer (2005), Mitman and Rabinovich (2015), and others. Following Setty and Yedid-Levi (2021), we calibrate the maximum payout $\chi$ to be 0.8433, which amounts to 48% of the median wage in the model. The eligibility probability $P_{e}$ is chosen to be 0.5385 to generate an average unemployment benefit duration of 26 weeks, as in Mitman and Rabinovich (2015).
@@ 3.5 Labor search and wage bargain
The labor market turnover statistics are obtained from Cajner et al. (2017). The elasticity of the matching function, $\iota$, targets the job-finding rate of the black workers, and our calibrated value of 1.3012 is close to that of 1.25 in Den Haan et al. (2000) and Petrosky-Nadeau et al. (2018). We choose firm-type specific job destruction shocks $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers, respectively. The ensuing values for $\lambda_{np}$ and $\lambda_{p}$ are 0.0644 and 0.0268, respectively. [^6] All the black workers work only with $np$ firms, and hence $\lambda_{np}$ is exactly equal to the job separation rate of the black workers. On the other hand, white workers work in both $p$ and $np$ firms; hence, $\lambda_{p}$ (0.0268) is smaller than the aggregate separation rate of the white workers (0.0380). The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job-finding rate of white workers, while the posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to target the aggregate labor market tightness, $\theta$, of 1.0, following Wolcott (2021). [^7] Aggregate labor market tightness $\theta=\frac{v_{np}+v_{p}}{u_{bl}+u_{wh}}$. We find that the prejudiced firms pay a lot more than non-prejudiced firms to post their vacancies, with $\kappa_{p}$ calibrated to be 4.2622, compared to $\kappa_{np}$ taking a value of 2.735. The bargaining power of black workers, $\xi_{bl}$, is calibrated to target the average black-white racial wage ratio of 0.75 (Derenoncourt and Montialoux, 2021). The white worker’s bargaining power, $\xi_{wh}$, is chosen to generate the average firm profit share of 3.3% as in Nakajima (2012). Consistent with our expectations, we find that the black workers have a lower bargaining power compared to the white workers, with $\xi_{bl}$ to be 0.1397 and $\xi_{wh}$ to be 0.2110.
@@ 3.6 Assessing the model as a quantitative theory of racial disparity
Our calibration successfully captures the racial gaps across three important dimensions: income, wealth, and employment outcomes. We now discuss these dimensions in more detail before proceeding with the quantitative exercises that examine the macroeconomic impact of racial discrimination.
@@
Note: This figure compares the model-generated steady-state distributions of wealth and labor income with their empirical analogs. The empirical labor income and wealth distributions are estimated from the 2007 American Community Survey (ACS) and the 2007 Consumer Finance Survey (SCF).
First, the model captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we do not target these distributions explicitly. As with most incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile. On the other hand, we are much closer to the empirical distributions in the lower quintiles since we target the share of zero-wealth workers in our calibration.
@@
Second, the model reproduces the disparities in unemployment, labor income, and wealth between black and white workers. As targeted in the calibration, the model replicates the empirical racial labor income gap and the empirical racial difference in the share of people at zero wealth. In addition to the targeted moments, Table 2 demonstrates that our model successfully generates the untargeted employment and wealth moments. Specifically, our model produces a 5% unemployment rate among white workers compared to 12% among black workers. Apart from the lower separation rate, white workers have a lower unemployment rate due to their access to prejudiced firms. On the wealth dimension, without targeting, the model generates a mean black-white wealth ratio of 0.28, close to 0.23 as in Kaplan et al. (2014). However, the model understates the median wealth gap between black and white workers. The median black worker holds 33% of the wealth of the median white worker in our model compared to 17% in Kaplan et al. (2014).
In summary, the model reflects the racial disparities present in labor income, wealth, and unemployment outcomes. Next, we will investigate how racial discrimination in hiring contributes to these disparities.
@@ 4.1 Sustaining racial discrimination
The labor search-and-matching framework allows firms to earn a profit from the surplus generated by search frictions. This feature permits prejudiced and non-prejudiced firms to coexist in a competitive equilibrium. It contrasts with a canonical frictionless neoclassical model where prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers as described in Becker (1957).
Table 3 shows firm performances in the steady state. $p$ firms have a total profit of 0.04, while $np$ firms have 0.03. This difference arises because $p$ firms experience a lower job destruction shock (0.0268) than $np$ firms (0.0644). Although $p$ firms incur higher vacancy posting costs (4.2622) to hire only white workers, they retain workers longer (employment level at 0.54) at lower posting rates (0.02). Conversely, $np$ firms face lower vacancy posting costs (2.735), but their high job destruction shock leads to a lower employment level (0.40) and a more frequent need for vacancy postings (0.05).
Table 3: Firms in the steady state
Moments Benchmark
$p$ firm profit 0.04
np firm profit 0.03
p firm vacancy 0.02
np firm vacancy 0.05
p firm employment 0.54
np firm employment 0.40
Note: This table shows steady state profit, vacancy postings, and employment levels for $p$ and $np$ firms in the benchmark model.
Next, we examine the steady-state interplay between prejudiced and non-prejudiced firms. At the benchmark equilibrium, $p$ firms post a positive number of vacancies with sufficient expected profits to cover the posting costs. We gradually increase penalties for them by raising $\kappa_{p}$ until their expected operating profit no longer justifies vacancy postings. Figure 2 displays the profit, vacancy, and employment dynamics between the $p$ and $np$ firms. As $\kappa_{p}$ rises, $p$ firms become increasingly disadvantaged, gradually losing their profits, vacancy postings, and employment. Conversely, with fewer $p$ firms competing in the market, $np$ firms see a steady increase in profits, vacancy postings, and employment.
It is worth noting that the dynamics of $p$ and $np$ firms are not symmetric, given the differences in vacancy posting costs and job destruction shocks. As $p$ firms’ profits decline, the profits of $np$ firms increase at a slower rate. Additionally, as $np$ firms replace $p$ firms, $np$ firms recruit from both black and white workers, rather than only white workers displaced from $p$ firms. This is mainly reflected in $np$ firms’ more elastic vacancy posting response.
In the limit, as $\kappa_{p}$ approaches the total expected value of an operating $p$ firm, posting vacancies exclusively for white workers becomes entirely unprofitable. As a result, all $p$ firms exit the market. The value of $\kappa_{p}$ at which $p$ firms exit entirely is 217% of the benchmark value of 4.2622. Meanwhile, $np$ firms become the sole employers in the economy. The gray dotted lines in Figure 2 illustrate a steady state where we construct a counterfactual model that structurally removes $p$ firms from the benchmark model. Thus, increasing $\kappa_{p}$ in the limit asymptotically equals the alternative of structurally eliminating $p$ firms.
@@ 4.2 Impact of hiring discrimination on racial inequality
In this subsection, we compare the benchmark economy to those without prejudiced firms to illustrate the aggregate and distributional impact of discriminatory hiring. Table 4 presents the comparisons. All firms in the counterfactual models provide equal hiring opportunities to black and white workers. Column 2 removes the $p$ firms from the benchmark model, which leaves all workers with equal job-finding rates with the $np$ firms and equal job destruction shocks at the benchmark $\lambda_{np}$. In Column 3, we further adjust the job destruction shocks to be race-specific, with $\lambda_{wh}$ and $\lambda_{bl}$ designating the shocks for white and black workers, respectively. [^8] In the benchmark economy, we structure firm-specific job destruction shocks ($\lambda_{np}$ and $\lambda_{p}$) to match race-specific job separation rates. This counterfactual economy only has one type of firm. We assign white workers a job destruction shock, $\lambda_{wh}$, of 0.0380 and black workers a destruction shock, $\lambda_{bl}$, of 0.0644 to maintain black and white workers’ job separation rates as in the calibration. Reassigning the job destruction shocks further isolates the impact of job separation.
@@
The first panel illustrates the changes in racial employment, wage, and wealth disparities after removing prejudiced firms. The first two rows describe the job separation rates. By construction, Column 2 equalizes job separation rates by race, while Column 3 retains the differences as the benchmark calibration. The next two rows describe the job-finding rates. Since there are no $p$ firms in the economy in Columns 2 and 3, the job-finding rates are equal for black and white workers. White workers have only one sector for employment; hence, their job-finding rate decreases. Black workers have more job opportunities as the $np$ sector posts more vacancies; thus, their job-finding rate increases. Comparing Columns 2 and 3, the much lower job destruction shock for white workers in Column 3 results in fewer workers searching for jobs, which leads to higher job-finding rates than in Column 2. Altogether, removing $p$ firms reduces the racial gap in the unemployment rate. In Column 2, the racial unemployment rate gap completely closes as job turnover equalizes. In Column 3, the unemployment rate gap closes from 7% to 3% due to the remaining racial difference in $\lambda$.
Turning to the wage impact, black workers can bargain for higher wage rates as they receive more job opportunities when we remove discriminatory hiring. White workers, however, no longer have access to the exclusive $p$ sector; their outside options are lower during bargaining, thus driving down their wage rates. As a result, the average black-to-white wage ratio increases from 75% in the benchmark to 88% in Column 2 and to 86% in Column 3. We conclude that removing discriminatory hiring closes 52% (Column 2) to 44% (Column 3) of the average wage gap between black and white workers.
Additionally, removing discriminatory hiring reduces the wealth gap, with the mean wealth ratio moving from 0.28 to 0.33 and the median wealth ratio increasing from 0.33 to 0.44. The results remain stable between Columns 2 and 3. We infer that discriminatory hiring explains about 7% of the mean wealth gap and 16% of the median wealth gap.
@@
The second panel displays the changes in firm attributes. [^10] We loosely refer to the one-sector economy without discriminatory hiring as $np$ firms, though the Column 3 specification differs from the benchmark $np$ firms in their job destruction shock. Without prejudiced firms, $np$ firms’ profits double, and vacancy postings and employment levels increase. These changes result in the aforementioned advantages for black workers in the counterfactual economies. In Column 3, the lower job destruction shock for white workers further drives up the overall $np$ employment and reduces vacancy postings compared to Column 2.
The third panel of Table 4 displays the labor market conditions. As $p$ firms exit, a greater number of $np$ firms enter the market endogenously. The labor market tightness for the $np$ sector, $\theta_{np}$, jumps from 0.73 in the benchmark to 1.02 in the equilibrium without $p$ firms, and to 1.12 when we further split $\lambda$ by race. The ensuing market tightness is higher than the benchmark aggregate market tightness of 1.
The last panel of Table 4 describes the aggregate outcomes. With more $np$ firms entering, the aggregate output remains relatively unchanged in Column 3. The capital-output ratio slightly decreases to 10.25. The economy-wide average wage increases to 2.05 from the benchmark of 2.04, and the aggregate unemployment rate decreases to 6%. However, Column 2 experiences a more significant reduction in output and average wage, along with a higher unemployment rate due to the overall higher job destruction shock that requires firms to frequently pay for vacancy posting costs instead of using resources for production.
Our results show that the presence of prejudiced firms can explain sizable fractions of income and wealth inequality. Penalizing prejudiced firms primarily improves employment opportunities and wage outcomes for black workers; it also worsens the situation for white workers. Removing $p$ firms decreases job-finding rates and increases unemployment for white workers, lowering their bargaining positions and affecting their income and wealth. To quantify the differential effects of removing discriminatory hiring, we next calculate the welfare changes experienced by black and white workers.